OpenStudy (anonymous):

Can somebody refresh my memory on how to do these? I want to say y=mx+b but i cant remeber which number goes where? What is the equation of the line that has a slope of -1 and goes through point (0, 5)?

OpenStudy (anonymous):

m = slope and b is the y-intercept. to write the equation of a line you first substitute the point and the slope into the equation y = mx+b so here it is... 5=-1(0) +b 5 = 0 + b 5=b so your equation is y = -x + 5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

sure